# Revisiting Isocurvature Bounds on the Minimal QCD Axion

### v1

@@ Abstract
The QCD axion has important connections to early universe cosmology. For example, it is often said that isocurvature limits rule out a combination of high axion decay constant, $f_{a}$, and high inflationary Hubble scale, $H_{I}$. High scales are theoretically motivated, so it is important to ask how robust this constraint is. We demonstrate that this constraint is naturally evaded when the quartic coupling of the complex $U(1)_{\textsc{pq}}$-breaking field is small. In this case, $f_{a}$ changes from a larger value during inflation to a smaller value in the later universe, suppressing isocurvature perturbations. Importantly, we show that in large parts of parameter space this solution is not jeopardised by overproduction of the axion through parametric resonance. The isocurvature bounds are thus dependent on UV physics. We have found that, even for the minimal QCD axion, large parts of UV parameter space at both high $f_{a}$ and high $H_{I}$ are in fact allowed, not ruled out by isocurvature constraints.
@@ 1.1 Background
The QCD axion is one of the best theoretically-motivated extensions to the Standard Model of particle physics (see e.g. [1, 2] for reviews). It solves the strong CP problem and also gives a natural dark matter candidate. The question of the production of axion dark matter then becomes a critical one. Interestingly, the axion dark matter abundance today is sensitive to initial conditions for the axion field arising from inflation at the beginning of the universe. This means that cosmological measurements of inflation may affect possibilities for axion dark matter and similarly any direct observations of axion dark matter could inform our understanding of the beginning of the universe.
@@
In the pre-inflationary regime, PQ symmetry is already broken before the end of inflation. In this case, we generically expect that the axion field will initially have a (roughly) homogeneous value, $\theta_{\text{i}}$, over the entire observable universe. Then in order to produce the observed dark matter abundance for a given $\theta_{\text{i}}$, a particular value for the axion decay constant $f_{a}$ (or equivalently the axion mass $m_{a}$) must be chosen which satisfies Eq. 2.7. This means that at least one of the two dimensionless constants of the theory $\theta_{\text{i}}$ and $\frac{f_{a}}{M_{\textsc{p}}}$ must be small. UV considerations motivate a large $f_{a}$ (see e.g. [3]), which would then require a small $\theta_{\text{i}}$. This small $\theta_{\text{i}}$ can occur naturally for the QCD axion with a long period of inflation (the ‘stochastic axion’ scenario) [4, 5], and additionally is explained naturally in many models (see e.g. [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]).
There are restrictions on the allowed parameter space for the pre-inflationary regime due to isocurvature constraints (see e.g. [17, 18, 19, 20, 21]). The value of the axion field will acquire quantum fluctuations during inflation. In the later universe these become fluctuations of the axion density which are uncorrelated with the regular adiabatic density perturbations, and hence this gives rise to isocurvature fluctuations. The region of parameter space which was claimed to be excluded by isocurvature constraints is shown in Fig. 2. Note that it is possible to build models to evade these isocurvature constraints (see e.g. [22, 7, 23, 24, 25, 26, 27, 28, 29]). However the minimal model for the QCD axion plus inflation, which is what we consider here, is claimed to be constrained by isocurvature. As is apparent in the figure, larger Hubble scales during inflation are generally claimed to be excluded by these constraints for the pre-inflationary axion scenario.
@@ 1.2 Executive Summary
In order for this mechanism to work, there must be one small parameter in the UV Lagrangian of the saxion. The ratio of the saxion quartic $\lambda_{\Phi}$ to its gravitational coupling $\xi$ must be small (though note that $\xi$ can just be $\mathcal{O}(1)$). The smaller this ratio is, the larger a region can be rescued from isocurvature constraints. There are known models which naturally have a small effective quartic or very flat potential for the saxion, e.g. in supergravity [30, 31, 32, 33]. The kinetic (or acoustic) misalignment mechanism tends to rely on this as well in order to enhance the energy density associated to the axion rotation (see e.g. [34, 35, 36, 37, 38, 39, 40, 41]). Kinetic misalignment is also similar in often using a negative, Hubble-induced mass during inflation which drives the saxion to large field values (giving a large $f_{a}$ during inflation). Other similar mechanisms have been proposed with a flat potential and a large initial saxion field value, for example to generate the baryon asymmetry from the axion [42, 43].
The possibility of reducing isocurvature constraints through such a mechanism has previously been considered [44, 45, 46, 47, 48, 25, 49, 34, 36, 27]. However it was worried that parametric resonance may generate too large an axion abundance, ruling out this scenario. We demonstrate that in fact there is a significant amount of viable parameter space where isocurvature bounds are removed and yet parametric resonance is not strong enough to overproduce the axion. As we show, the strength of the parametric resonance depends on many detailed aspects of the UV scenario, including the model of inflation and reheating and the saxion potential. Such UV parameters are hidden on the simple plots of $H_{I}$ vs $f_{a}$ as in Fig. 1, but in fact determine whether the axion model will be ruled out by isocurvature. For example, in different models of inflation and reheating the transition in PQ breaking scale from the high scale $f_{a}^{\text{(inf)}}$ down to the scale in the late universe $f_{a}$ can occur at different rates. The more adiabatic this transition is, the weaker the parametric resonance is. We investigate a couple different inflation models and display the allowed parts of parameter space in Fig. 6. Considering other inflation/reheating models would likely expand the allowed region even further.
@@ 2 Standard isocurvature constraints on axion dark matter
In this section, we concisely review the standard constraints on axion dark matter (DM) from isocurvature perturbations. According to CMB measurements, matter curvature perturbations spatially track the perturbations of SM radiation $\gamma$ around the time of matter-radiation equality. Any deviation for a species $a$ is parametrised by isocurvature perturbations $\mathcal{S}_{a\gamma}(\mathbf{x})=3(\zeta_{a}(\mathbf{x})-\zeta_{\gamma}(\mathbf{x}))$, where $\zeta_{i}(\mathbf{x})=\Psi(\mathbf{x})-H\delta\rho_{i}(\mathbf{x})/\dot{\rho}_{i}$ is the curvature perturbation of the species $i=a,\gamma$ [50]. The strongest observational bound [^1] Other interesting signatures of dark matter isocurvature may appear in correlators of large-scale structures [21]. on isocurvature perturbations on large scales comes from the CMB measurements of Planck18 [51, Table 14]. [^2] The very recent results by ACT [52, Table 6] are mildly weaker, with Planck18 + ACT giving $\beta_{\text{iso}}<0.042$, and Planck18 + ACT + DESI 1y BAO yielding $\beta_{\text{iso}}<0.049$. The limit on the power spectrum $\mathcal{P}_{a}$ of isocurvature perturbations $\mathcal{S}_{a\gamma}(\mathbf{x})$ that is relevant to our discussion is
$\mathcal{P}_{a}<A_{s}\beta_{\text{iso}}\,,\quad A_{s}=2.1\cdot 10^{-9},\quad\quad\beta_{\text{iso}}=0.038\,,$ (2.1)
@@
We now specialise the discussion to isocurvature perturbations of axion DM, which we can compute at a time during the early period of radiation domination when the axion has started oscillating around its minimum, so that its equation of state is that of non-relativistic matter, $\rho_{a}(\mathbf{x})=m\,n_{a}(\mathbf{x})$ and $\dot{\rho}_{a}=-3H\rho_{a}$. We consider the spatially flat gauge $\Psi=0$. At early times, deep in radiation domination, the axion is a tiny component of the total energy density, and $\delta\rho_{a}/\rho_{a}\gg\delta\rho_{a}/\rho_{\gamma}\sim\delta\rho_{\gamma}/\rho_{\gamma}$, so that we can approximate $\mathcal{S}_{a\gamma}=3\zeta_{a}$. We can then write the contribution from the axion field to $\mathcal{P}_{a}$ as [19, 20]
$\mathcal{P}_{a}=\left\langle\left(\frac{\delta n_{a}}{n_{a}}\right)^{2}\right\rangle,\quad n_{a}(\mathbf{x})=\frac{\rho_{a}(\mathbf{x})}{m_{a}}=\frac{1}{2}m_{a}f_{a}^{2}\theta(\mathbf{x})^{2}\,\mathscr{F}_{\text{anh.}}\big{(}\theta(\mathbf{x})\big{)}$ (2.2)
where $\theta(x)\equiv a(x)/f_{a}$, while $\mathscr{F}_{\text{anh.}}(\theta)$ accounts for anharmonicity effects and differs from 1 only for $\theta\gtrsim 2$. In this work, we are interested in a region of parameter space where $\theta<1$, so we ignore anharmonic effects for simplicity (although we show the correct isocurvature curve in our plots including those effects). We define the average value $\theta_{\text{i}}\equiv\langle\theta(x)\rangle$ for the initial misalignment angle in our observable Universe, and the spatial fluctuations $\delta\theta(\mathbf{x})\equiv\theta(\mathbf{x})-\theta_{\text{i}}$. We assume that $\delta\theta$ is a Gaussian variable, with standard deviation set by quantum fluctuations during inflation
$\sigma_{\theta}\equiv\sqrt{\langle\delta\theta(\mathbf{x})^{2}\rangle}=\frac{H_{I}}{2\pi f_{a}^{\text{(inf)}}}\sqrt{N_{\textsc{cmb}}}\,.$ (2.3)
Here, $f_{a}^{\text{(inf)}}$ is the vev of the radial direction of the complex PQ field (which we dub the saxion) during inflation, that we allow to differ from its present-day value $f_{a}$. The parameter $N_{\textsc{cmb}}\gtrsim 1$ quantifies the growth of the variance of the axion field during the range of $e$-folds contributing to the experimental constraints on isocurvature, and we take $N_{\textsc{cmb}}=1$ for simplicity. [^3] Previous analyses provided arguments for e.g. $\sqrt{N_{\textsc{cmb}}}=2$ [19] or 4 [17].
@@
$\frac{\delta n_{a}}{n_{a}}\approx\frac{\theta^{2}-\langle\theta^{2}\rangle}{\langle\theta^{2}\rangle}=\frac{\big{(}\theta_{\text{i}}^{2}+2\theta_{\text{i}}\delta\theta+(\delta\theta)^{2}\big{)}-\big{(}\theta_{\text{i}}^{2}+\sigma_{\theta}^{2}\big{)}}{\theta_{\text{i}}^{2}+\sigma_{\theta}^{2}}=\frac{2\theta_{\text{i}}\delta\theta+(\delta\theta)^{2}-\sigma_{\theta}^{2}}{\theta_{\text{i}}^{2}+\sigma_{\theta}^{2}}\,.$ (2.4)
@@
$\left\langle\left(\frac{\delta n_{a}}{n_{a}}\right)^{2}\right\rangle\approx\frac{\langle\delta\theta^{4}+4\theta_{\text{i}}^{2}(\delta\theta)^{2}+\sigma_{\theta}^{4}-2(\delta\theta)^{2}\sigma_{\theta}^{2}+(\text{odd in }\delta\theta)\rangle}{(\theta_{\text{i}}^{2}+\sigma_{\theta}^{2})^{2}}=\frac{4\theta_{\text{i}}^{2}\sigma_{\theta}^{2}+2\sigma_{\theta}^{4}}{(\theta_{\text{i}}^{2}+\sigma_{\theta}^{2})^{2}}\overset{\theta_{\text{i}}\gg\sigma_{\theta}}{\longrightarrow}\frac{4\sigma_{\theta}^{2}}{\theta_{\text{i}}^{2}}\,,$ (2.5)
@@
$\frac{H_{I}}{f_{a}^{\text{(inf)}}\theta_{\text{i}}}<\sqrt{\frac{\pi^{2}A_{s}\beta_{\text{iso}}}{N_{\textsc{cmb}}}}=2.8\cdot 10^{-5}\,.$ (2.6)
The initial misalignment $\theta_{\text{i}}$ yielding the DM abundance for the axion field is in one-to-one correspondence with the axion decay constant $f_{a}$, and encodes various features of the axion physics at the epoch of the QCD phase transition. In the dilute instanton gas approximation for the axion potential (approximated to a quadratic potential for $\theta_{\text{i}}\lesssim 2$), with the simplification that the axion starts oscillating exactly at $m_{a}=3H(T_{\text{osc}})$, assuming that the temperature-dependent axion mass scales as $m_{a}(T)=0.026\,m_{a}(160\,\text{MeV}/T)^{4}$, one finds [53, 54]
$\theta_{\text{i}}=\left(\frac{7.4\cdot 10^{11}\,\text{GeV}}{f_{a}}\right)^{7/12}$ (2.7)
@@
This initial misalignment angle $\theta_{\text{i}}$ for the QCD axion to constitute DM is marked on the left axis in Fig. 2. The combination of Eqs. 2.6 and 2.7, together with the identification $f_{a}^{\text{(inf)}}=f_{a}$, defines the bound from isocurvature plotted with a grey line in the parameter space $(H_{I},f_{a})$ of Fig. 2. The separation between the pre-inflationary regime (where $U(1)_{\textsc{pq}}$ is not restored after inflation, and the misalignment angle $\theta_{\text{i}}$ in our observable Universe is set before inflation) and the post-inflationary regime (with $U(1)_{\textsc{pq}}$ restored after inflation) depends on the highest temperature $T_{\text{max}}$ reached after inflation. [^4] Strictly speaking, by $T_{\text{max}}$ we refer to the largest temperature reached by species that contribute in a sizeable amount to the thermal effective potential of the saxion. If the reheating of the SM and of the sector charged under $U(1)_{\textsc{pq}}$ occurs later, the relevant $T$ for Eq. 2.8 can be smaller than $T_{\text{max}}$. The maximum possible temperature $T_{\mathrm{inst.\,RH}}$ corresponds to an instantaneous reheating, $\tfrac{\pi^{2}}{30}g_{\star}(T_{\mathrm{inst.\,RH}})^{4}=3H_{e}^{2}M_{\textsc{p}}^{2}$, with $H_{e}$ the Hubble rate at the end of inflation (and generally $H_{e}\leq H_{I}$), and we set $g_{\star}=106.75$ for definiteness. The condition to restore $U(1)_{\textsc{pq}}$ after inflation and thus to be in the post-inflation regime is then [^5] Some references quote $H_{I}/(2\pi)$ rather than $H_{I}$, which has a pretty small quantitative impact on Eq. 2.8. [4, 55, 56, 57]
[…]
which is of course dependent on the reheating temperature. We show in Fig. 2 the boundary corresponding to $T_{\text{max}}=0.1\,T_{\mathrm{inst.\,RH}}$ for reference. The two upper boundaries on $H_{I}$ and $f_{a}$ are given respectively by the upper limit $r<0.036$ (or $H_{I}<4.6\cdot 10^{13}\,\text{GeV}$) from BICEP-Planck [58] and a selection of bounds from black-hole superradiance [59, 60, 61, 62, 63, 64, 65, 66, 67].

### v2

@@ Abstract
The QCD axion has important connections to early universe cosmology. For example, it is often said that isocurvature limits rule out a combination of high axion decay constant, $f_{a}$, and high inflationary Hubble scale, $H_{I}$. High scales are theoretically motivated, so it is important to ask how robust this constraint is. We demonstrate that this constraint is naturally evaded when the quartic coupling of the complex $U(1)_{\textsc{pq}}$-breaking field is small (e.g. $\lesssim 10^{-6}$). In this case, $f_{a}$ changes from a larger value during inflation to a smaller value in the later universe, suppressing isocurvature perturbations. Importantly, we show that in large parts of parameter space this solution is not jeopardised by overproduction of the axion through parametric resonance. The isocurvature bounds are thus dependent on UV physics. We have found that, even for the minimal QCD axion, large parts of UV parameter space at both high $f_{a}$ and high $H_{I}$ are in fact allowed, not ruled out by isocurvature constraints.
@@ 1.1 Background
The QCD axion is one of the best theoretically-motivated extensions to the Standard Model of particle physics (see e.g. Graham:2015ouw; Baryakhtar:2025jwh for reviews). It solves the strong CP problem and also gives a natural dark matter candidate. The question of the production of axion dark matter then becomes a critical one. Interestingly, the axion dark matter abundance today is sensitive to initial conditions for the axion field arising from inflation at the beginning of the universe. This means that cosmological measurements of inflation may affect possibilities for axion dark matter and similarly any direct observations of axion dark matter could inform our understanding of the beginning of the universe.
@@
In the pre-inflationary regime, PQ symmetry is already broken before the end of inflation. In this case, we generically expect that the axion field will initially have a (roughly) homogeneous value, $\theta_{\text{i}}$, over the entire observable universe. Then in order to produce the observed dark matter abundance for a given $\theta_{\text{i}}$, a particular value for the axion decay constant $f_{a}$ (or equivalently the axion mass $m_{a}$) must be chosen which satisfies Eq. 7. This means that at least one of the two dimensionless constants of the theory $\theta_{\text{i}}$ and $\frac{f_{a}}{M_{\textsc{p}}}$ must be small. UV arguments motivate considering a wide range for $f_{a}$ (see e.g. Svrcek:2006yi; Demirtas:2020dbm; Gendler:2023kjt; Gendler:2024adn; Petrossian-Byrne:2025mto; Loladze:2025uvf; Benabou:2025kgx), including large values which would then require a small $\theta_{\text{i}}$. This small $\theta_{\text{i}}$ can occur naturally for the QCD axion with a long period of inflation (the ‘stochastic axion’ scenario) Graham:2018jyp; Takahashi:2018tdu, and additionally is explained naturally in many models (see e.g. Agrawal:2017eqm; Nomura:2015xil; Dine:1982ah; Steinhardt:1983ia; Lazarides:1990xp; Kawasaki:1995vt; Dvali:1995ce; Choi:1996fs; Banks:1996ea; Banks:2002sd; Bao:2022hsg).
There are restrictions on the allowed parameter space for the pre-inflationary regime due to isocurvature constraints (see e.g. Lyth:1991ub; Fox:2004kb; Hertzberg:2008wr; Visinelli:2009zm; Chen:2023txq). The value of the axion field will acquire quantum fluctuations during inflation. In the later universe these become fluctuations of the axion density which are uncorrelated with the regular adiabatic density perturbations, and hence this gives rise to isocurvature fluctuations. The region of parameter space which was claimed to be excluded by isocurvature constraints is shown in Fig. 2. Note that it is possible to build models to evade these isocurvature constraints (see e.g. Jeong:2013xta; Nomura:2015xil; Kawasaki:2015lpf; Takahashi:2015waa; Harigaya:2015hha; Nakayama:2015pba; Ballesteros:2021bee; Choi:2014uaa; Kawasaki:2014una). However the minimal model for the QCD axion plus inflation, which is what we consider here, is claimed to be constrained by isocurvature. As is apparent in the figure, larger Hubble scales during inflation are generally claimed to be excluded by these constraints for the pre-inflationary axion scenario.
@@ 1.2 Executive Summary
In order for this mechanism to work, there must be one small parameter in the UV Lagrangian of the saxion. The ratio of the saxion quartic $\lambda_{\Phi}$ to its gravitational coupling $\xi$ must be small (though note that $\xi$ can be $\mathcal{O}(1)$). The exact values of these parameters are dependent on the UV models of inflation and the saxion. For the particular example we consider in most detail, the quartic must be $\lambda_{\Phi}\lesssim 10^{-6}$ in order to start cutting into the isocurvature bound. The smaller this ratio $\lambda_{\Phi}/\xi$ is, the larger a region can be rescued from isocurvature constraints (for reasons discussed below we choose to cut off our scan at a lower limit of $\lambda_{\Phi}=10^{-18}$ though one could have continued lower). There are known models which naturally have a small effective quartic or very flat potential for the saxion, e.g. in supergravity Moxhay:1984am; Abe:2001cg; Nakamura:2008ey; DEramo:2015iqd. The kinetic (or acoustic) misalignment mechanism tends to rely on this as well in order to enhance the energy density associated to the axion rotation (see e.g. Co:2017mop; Co:2019jts; Co:2020dya; Domcke:2022wpb; Eroncel:2022vjg; Eroncel:2024rpe; Bodas:2025eca; Eroncel:2025qlk). Kinetic misalignment is also similar in often using a negative, Hubble-induced mass during inflation which drives the saxion to large field values (giving a large $f_{a}$ during inflation). Other similar mechanisms have been proposed with a flat potential and a large initial saxion field value, for example to generate the baryon asymmetry from the axion Co:2019wyp; Co:2023mhe.
The possibility of reducing isocurvature constraints through such a mechanism has previously been considered Linde:1991km; Kasuya:1996ns; Kawasaki:2008jc; Folkerts:2013tua; Fairbairn:2014zta; Kearney:2016vqw; Kawasaki:2013iha; Harigaya:2015hha; Kobayashi:2016qld; Co:2017mop; Co:2020dya; Ballesteros:2021bee. However it was worried that parametric resonance may generate too large an axion abundance, ruling out this scenario. We demonstrate that in fact there is a significant amount of viable parameter space where isocurvature bounds are removed and yet parametric resonance is not strong enough to overproduce the axion. As we show, the strength of the parametric resonance depends on many detailed aspects of the UV scenario, including the model of inflation and reheating and the saxion potential. Such UV parameters are hidden on the simple plots of $H_{I}$ vs $f_{a}$ as in Fig. 1, but in fact determine whether the axion model will be ruled out by isocurvature. For example, in different models of inflation and reheating the transition in PQ breaking scale from the high scale $f_{a}^{\text{(inf)}}$ down to the scale in the late universe $f_{a}$ can occur at different rates. The more adiabatic this transition is, the weaker the parametric resonance is. We investigate a couple different inflation models and display the allowed parts of parameter space in Fig. 6. Considering other inflation/reheating models would likely expand the allowed region even further.
@@ 2 Standard isocurvature constraints on axion dark matter
In this section, we concisely review the standard constraints on axion dark matter (DM) from isocurvature perturbations. According to CMB measurements, matter curvature perturbations spatially track the perturbations of SM radiation $\gamma$ around the time of matter-radiation equality. Any deviation for a species $a$ is parametrised by isocurvature perturbations $\mathcal{S}_{a\gamma}(\mathbf{x})=3(\zeta_{a}(\mathbf{x})-\zeta_{\gamma}(\mathbf{x}))$, where $\zeta_{i}(\mathbf{x})=\Psi(\mathbf{x})-H\delta\rho_{i}(\mathbf{x})/\dot{\rho}_{i}$ is the curvature perturbation of the species $i=a,\gamma$ Kodama:1984ziu. The strongest observational bound [^1] Other interesting signatures of dark matter isocurvature may appear in correlators of large-scale structures Chen:2023txq. on isocurvature perturbations on large scales comes from the CMB measurements of Planck18 (Planck:2018jri, Table 14). [^2] The very recent results by ACT (ACT:2025tim, Table 6) are mildly weaker, with Planck18 + ACT giving $\beta_{\text{iso}}<0.042$, and Planck18 + ACT + DESI 1y BAO yielding $\beta_{\text{iso}}<0.049$. The limit on the power spectrum $\mathcal{P}_{a}$ of isocurvature perturbations $\mathcal{S}_{a\gamma}(\mathbf{x})$ that is relevant to our discussion is
$\mathcal{P}_{a}<A_{s}\beta_{\text{iso}}\,,\quad A_{s}=2.1\cdot 10^{-9},\quad\quad\beta_{\text{iso}}=0.038\,$, (1)
@@
We now specialise the discussion to isocurvature perturbations of axion DM, which we can compute at a time during the early period of radiation domination when the axion has started oscillating around its minimum, so that its equation of state is that of non-relativistic matter, $\rho_{a}(\mathbf{x})=m\,n_{a}(\mathbf{x})$ and $\dot{\rho}_{a}=-3H\rho_{a}$. We consider the spatially flat gauge $\Psi=0$. At early times, deep in radiation domination, the axion is a tiny component of the total energy density, and $\delta\rho_{a}/\rho_{a}\gg\delta\rho_{a}/\rho_{\gamma}\sim\delta\rho_{\gamma}/\rho_{\gamma}$, so that we can approximate $\mathcal{S}_{a\gamma}=3\zeta_{a}$. We can then write the contribution from the axion field to $\mathcal{P}_{a}$ as Hertzberg:2008wr; Visinelli:2009zm
$\mathcal{P}_{a}=\left\langle\left(\frac{\delta n_{a}}{n_{a}}\right)^{2}\right\rangle,\quad n_{a}(\mathbf{x})=\frac{\rho_{a}(\mathbf{x})}{m_{a}}=\frac{1}{2}m_{a}f_{a}^{2}\theta(\mathbf{x})^{2}\,\mathscr{F}_{\text{anh.}}\big(\theta(\mathbf{x})\big)$ (2)
where $\theta(x)\equiv a(x)/f_{a}$, while $\mathscr{F}_{\text{anh.}}(\theta)$ accounts for anharmonicity effects and differs from 1 only for $\theta\gtrsim 2$. In this work, we are interested in a region of parameter space where $\theta<1$, so we ignore anharmonic effects for simplicity (although we show the correct isocurvature curve in our plots including those effects, see e.g. Kobayashi:2013nva). We define the average value $\theta_{\text{i}}\equiv\langle\theta(x)\rangle$ for the initial misalignment angle in our observable Universe, and the spatial fluctuations $\delta\theta(\mathbf{x})\equiv\theta(\mathbf{x})-\theta_{\text{i}}$. We assume that $\delta\theta$ is a Gaussian variable, with standard deviation set by quantum fluctuations during inflation
$\sigma_{\theta}\equiv\sqrt{\langle\delta\theta(\mathbf{x})^{2}\rangle}=\frac{H_{I}}{2\pi f_{a}^{\text{(inf)}}}\sqrt{N_{\textsc{cmb}}}\,.$ (3)
Here, $f_{a}^{\text{(inf)}}$ is the vev of the radial direction of the complex PQ field (which we dub the saxion) during inflation, that we allow to differ from its present-day value $f_{a}$. The parameter $N_{\textsc{cmb}}\gtrsim 1$ quantifies the growth of the variance of the axion field during the range of $e$-folds contributing to the experimental constraints on isocurvature, and we take $N_{\textsc{cmb}}=1$ for simplicity. [^3] Previous analyses provided arguments for e.g. $\sqrt{N_{\textsc{cmb}}}=2$ Hertzberg:2008wr or 4 Lyth:1991ub.
@@
$\frac{\delta n_{a}}{n_{a}}\approx\frac{\theta^{2}-\langle\theta^{2}\rangle}{\langle\theta^{2}\rangle}=\frac{\big(\theta_{\text{i}}^{2}+2\theta_{\text{i}}\delta\theta+(\delta\theta)^{2}\big)-\big(\theta_{\text{i}}^{2}+\sigma_{\theta}^{2}\big)}{\theta_{\text{i}}^{2}+\sigma_{\theta}^{2}}=\frac{2\theta_{\text{i}}\delta\theta+(\delta\theta)^{2}-\sigma_{\theta}^{2}}{\theta_{\text{i}}^{2}+\sigma_{\theta}^{2}}\,.$ (4)
@@
$\left\langle\left(\frac{\delta n_{a}}{n_{a}}\right)^{2}\right\rangle\approx\frac{\langle\delta\theta^{4}+4\theta_{\text{i}}^{2}(\delta\theta)^{2}+\sigma_{\theta}^{4}-2(\delta\theta)^{2}\sigma_{\theta}^{2}+(\text{odd in }\delta\theta)\rangle}{(\theta_{\text{i}}^{2}+\sigma_{\theta}^{2})^{2}}=\frac{4\theta_{\text{i}}^{2}\sigma_{\theta}^{2}+2\sigma_{\theta}^{4}}{(\theta_{\text{i}}^{2}+\sigma_{\theta}^{2})^{2}}\overset{\theta_{\text{i}}\gg\sigma_{\theta}}{\longrightarrow}\frac{4\sigma_{\theta}^{2}}{\theta_{\text{i}}^{2}}\,$, (5)
@@
$\frac{H_{I}}{f_{a}^{\text{(inf)}}\theta_{\text{i}}}<\sqrt{\frac{\pi^{2}A_{s}\beta_{\text{iso}}}{N_{\textsc{cmb}}}}=2.8\cdot 10^{-5}\,.$ (6)
The initial misalignment $\theta_{\text{i}}$ yielding the DM abundance for the axion field is in one-to-one correspondence with the axion decay constant $f_{a}$, and encodes various features of the axion physics at the epoch of the QCD phase transition. In the dilute instanton gas approximation for the axion potential (approximated to a quadratic potential for $\theta_{\text{i}}\lesssim 2$), with the simplification that the axion starts oscillating exactly at $m_{a}=3H(T_{\text{osc}})$, assuming that the temperature-dependent axion mass scales as $m_{a}(T)=0.026\,m_{a}(160\,\text{MeV}/T)^{4}$, one finds GrillidiCortona:2015jxo; DiLuzio:2020wdo
$\theta_{\text{i}}=\left(\frac{7.4\cdot 10^{11}\,\text{GeV}}{f_{a}}\right)^{7/12}$ (7)
@@
This initial misalignment angle $\theta_{\text{i}}$ for the QCD axion to constitute DM is marked on the left axis in Fig. 2. The combination of Eqs. 6 and 7, together with the identification $f_{a}^{\text{(inf)}}=f_{a}$, defines the bound from isocurvature plotted with a grey line in the parameter space $(H_{I},f_{a})$ of Fig. 2. The separation between the pre-inflationary regime (where $U(1)_{\textsc{pq}}$ is not restored after inflation, and the misalignment angle $\theta_{\text{i}}$ in our observable Universe is set before inflation) and the post-inflationary regime (with $U(1)_{\textsc{pq}}$ restored after inflation) depends on the highest temperature $T_{\text{max}}$ reached after inflation. [^4] Strictly speaking, by $T_{\text{max}}$ we refer to the largest temperature reached by species that contribute in a sizeable amount to the thermal effective potential of the saxion. If the reheating of the SM and of the sector charged under $U(1)_{\textsc{pq}}$ occurs later, the relevant $T$ for Eq. 8 can be smaller than $T_{\text{max}}$. The maximum possible temperature $T_{\mathrm{inst.\,RH}}$ corresponds to an instantaneous reheating, $\tfrac{\pi^{2}}{30}g_{\star}(T_{\mathrm{inst.\,RH}})^{4}=3H_{e}^{2}M_{\textsc{p}}^{2}$, with $H_{e}$ the Hubble rate at the end of inflation (and generally $H_{e}\leq H_{I}$), and we set $g_{\star}=106.75$ for definiteness. The condition to restore $U(1)_{\textsc{pq}}$ after inflation and thus to be in the post-inflation regime is then [^5] Some references quote $H_{I}/(2\pi)$ rather than $H_{I}$, which has a pretty small quantitative impact on Eq. 8. Graham:2018jyp; Gorghetto:2018myk; Buschmann:2019icd; Redi:2022llj
$f_{a}<\max\left(T_{\text{max}},H_{I}\right)\quad\text{({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}post-inflationary} regime)}$ (8)
which is of course dependent on the reheating temperature. We show in Fig. 2 the boundary corresponding to $T_{\text{max}}=0.1\,T_{\mathrm{inst.\,RH}}$ for reference. The two upper boundaries on $H_{I}$ and $f_{a}$ are given respectively by the upper limit $r<0.036$ (or $H_{I}<4.6\cdot 10^{13}\,\text{GeV}$) from BICEP-Planck BICEP:2021xfz and a selection of bounds from black-hole superradiance Yoshino:2014wwa; Brito:2015oca; Brito:2017wnc; Cardoso:2018tly; Tsukada:2018mbp; Palomba:2019vxe; Kaplan:2019ako; Sun:2019mqb; Baryakhtar:2020gao.